OpenStudy (abcdefghijklmq):

Need help quick WILL FAN+MEDAL Which two consecutive whole numbers is 643-16-04-00-00_files/i0140000.jpg between? a. 8 and 9 b. 5 and 6 c. 7 and 8 d. 6 and 7

OpenStudy (abcdefghijklmq):

\[\sqrt{45}\] this is what the it is in the question

OpenStudy (michele_laino):

hint: we can write this: \[36 < 45 < 49\]

OpenStudy (abcdefghijklmq):

what do you mean by that though? can you explain the formula to figuring out this kind of problem please

OpenStudy (michele_laino):

Please, make the square root of all sides of the above inequality

OpenStudy (michele_laino):

what do you get?

OpenStudy (abcdefghijklmq):

you get A= 64 and 81? B=25 and 36? C= 49 and 64? D= 36 and 49?

OpenStudy (michele_laino):

no, the square root, like this: \[\sqrt {36} < \sqrt {45} < \sqrt {49} \]

OpenStudy (abcdefghijklmq):

I don't get it

OpenStudy (michele_laino):

why?

OpenStudy (abcdefghijklmq):

how do you get the square root?

OpenStudy (michele_laino):

it is simple starting from this inequality: \[36 < 45 < 49\] I apply the square root operator so I get: \[\sqrt {36} < \sqrt {45} < \sqrt {49} \]

OpenStudy (abcdefghijklmq):

oh, would the answer be D?, because it is \[\sqrt{36} and \sqrt{49}\]

OpenStudy (michele_laino):

that's right!

OpenStudy (abcdefghijklmq):

okay! thank you so much!

OpenStudy (michele_laino):

thank you! :)